# Enabling Large-Scale and High-Precision Fluid Simulations on Near-Term Quantum Computers

### v1

@@ Abstract
Quantum computational fluid dynamics (QCFD) offers a promising alternative to classical computational fluid dynamics (CFD) by leveraging quantum algorithms for higher efficiency. This paper introduces a comprehensive QCFD method implemented on a superconducting quantum computer, demonstrating successful simulations of steady Poiseuille flow and unsteady acoustic wave propagation. The Poiseuille flow simulation achieved a relative error of less than $0.2\%$, and the unsteady acoustic wave simulation solved a 5043-dimension matrix, marking the largest simulation on a quantum computer to date. Our approach bridges quantum and classical computing, adapting to quantum hardware constraints and offering scalable solutions for large-scale CFD problems, which paves the way for practical applications of near-term quantum computers in computational science.
@@ 1 Introduction
Computational fluid dynamics (CFD) has played a pivotal role in numerous engineering and scientific disciplines, providing invaluable insights into complex flow behavior. However, as computational demands increase and the complexity of fluid dynamics problems intensifies, the limitations of classical computing architectures become increasingly apparent. Quantum computational fluid dynamics (QCFD) emerges as a promising alternative, offering the potential to overcome these limitations and unlock unprecedented computational power. At its core, QCFD employs quantum algorithms and quantum computing hardware to simulate fluid dynamics phenomena with higher efficiency. The specific methods of QCFD are diverse, including Lagrangian-type methods [1, 2], explicit Euler-type methods [3], implicit Euler-type methods [4, 5], and methods based on quantum neural networks [6]. Focusing on solving complicated fluid governing equations, the linear system of equations plays a fundamental role in most classical CFD methods. While solving linear systems is the most time-consuming part of these methods, quantum linear solvers (QLS) promise an exponential speedup since first proposed in Ref. [7]. Tab. 1 reports these developments on QLS.
@@
Despite these prominent advancements, the implementation of QCFD still faces challenges based on near-term quantum devices, i.e., noisy intermediate-scale quantum (NISQ) devices [14]. These aforementioned QLS are designed in the context of fault-tolerant quantum computation (FTQC) era [15, 16]. In a fault-tolerant quantum computer, redundant noisy physical qubits are employed to encode an ideal “logical qubit” without error [17]. However, only a few logical qubits can be demonstrated so far [18, 19], preventing most QLS in Tab. 1 from near-term practicability.
@@
Some efforts are conducted to enable computing fluid flows on quantum computers. One methodology is to emulate the flow field as a quantum state, utilizing the quantum state evolution to emulate the flow simulation, as shown in [20, 21]. Another methodology tries a variational approach to solve the equation, such as employing variational quantum linear solvers (VQLS) [22, 23, 24] in CFD, including solving potential and Stokes flows and the advection-diffusion equation [25, 26, 27]. However, current implementations are limited to small sizes due to the inherent limitations of near-term quantum computers, which are significantly affected by noise. Additionally, encoding the entire flow field into a quantum state and reading it out into classical data using full quantum tomography is highly inefficient. These noise and readout issues impede the scalability of existing methods to larger problems.
@@
In this paper, we propose a comprehensive method that enables large-scale and high-precision fluid simulation on near-term quantum computers. We implement it on a superconducting quantum computer, successfully simulating steady Poiseuille flow and unsteady acoustic wave propagation to showcase the ability of our method. In the Poiseuille flow simulation, we achieve less than $0.2\%$ relative error for showcasing the accuracy of a single step. In the unsteady acoustic wave propagation simulation, we successfully simulated 1681 grid points. Each time step solves a 5043-dimension matrix, which is the largest simulation on the quantum computer reported so far.
@@
To mitigate experimental error arising from noisy devices and approximation error, we introduce Iterative-QLS, by introducing iterations to enhance the precision rather than employing a single VQLS. This cycle aims to diminish residual error by achieving synthetic accuracy levels even lower than the physical accuracy. To overcome the problem size constraints imposed by intermediate-scale quantum processors, the subspace method is employed. This approach enables the hybridized utilization of quantum computing and classical computing, enabling simulation on a larger flow field. Furthermore, to enhance the efficiency of extracting classical information from quantum states, we have developed the sparse tomography technique. In summary, we present an integrated QCFD protocol capable of deriving high-precision solutions for problems of arbitrary size.
@@
We regard our algorithm as a bridge between quantum and classical computers. According to our findings, our method can adapt to the quantum computer’s capabilities, enabling collaboration between quantum and classical computers. Although the current solution uses only a limited number of qubits due to hardware constraints, advancements in quantum computing offer the possibility of solving larger subspaces by investing more quantum resources. Our approach is also a general-purpose quantum-enhanced linear solver method, applicable to other potential applications in computational science, including but not limited to FEM, FDM, and optimization problems. With successful, high-precision demonstrations on large-scale CFD problems, this paper marks a step towards scalable quantum-enhanced solutions for CFD and paves the way for the practical application of near-term quantum computers.
@@ 2.1 Basic concepts in quantum computing
The basic information unit in quantum information is qubit, which is analogous to a classical bit but with the ability to exist in a superposition of states. Mathematically, a single qubit state can be represented as a vector in 2-dimension Hilbert space:
@@ 2.2 The technical path of QCFD using QLS
In CFD, one must solve nonlinear partial differential equations, such as the Navier-Stokes equations. These equations can be represented as
$\partial{Q}/\partial{t}=\mathcal{N}(Q),$ (1)
where $\mathcal{N}(Q)$ represents nonlinear operators. The first step is discretization, which involves breaking down the equations into a form that can be handled computationally. Both time and space discretization schemes are employed, which can be either explicit or implicit. While explicit methods have strict stability requirements limiting the time step, implicit methods are preferred for their ability to handle larger time steps and provide numerical stability, particularly for stiff differential equations. Therefore, we use implicit methods, which are formulated as
@@
$\vec{Q}^{t+1}-\vec{Q}^{t}=\vec{R}(\vec{Q}^{t+1}),$ (3)
where $\vec{Q}^{t+1}\in\mathbb{R}^{n}$, and $n$ represents the dimension of the discretized space. Generally, this leads to a system of nonlinear equations in $\vec{Q}^{t+1}$, necessitating a linearization procedure.
@@ 2.3 Challenges of implementing CFD simulation on NISQ devices
Currently, several QCFD algorithms have been proposed, but there are still challenges when solving specific CFD problems on Noisy Intermediate-Scale Quantum (NISQ) chips. Here, we list and analyze these challenges.
Firstly, due to the impact of quantum noise, the quantum state produced by a quantum circuit executed on NISQ chips is prone to errors. This inaccuracy not only results in low-precision, unreliable solutions but also hinders the computation of unsteady flow using time-stepping methods.
Secondly, the number of qubits on NISQ chips is limited. This limitation arises not only because the number of qubits that can be fabricated on a quantum processor is restricted, but also because quantum noise limits the size of the circuit, preventing the number of qubits used in a quantum circuit from increasing [35].
@@ 3.1 Overview
When solving large-scale problems, the number of qubits required by quantum CFD algorithms may exceed the chip’s capacity. Therefore, a key issue for quantum CFD algorithms is how to solve large-scale CFD problems on NISQ devices with a limited number of qubits. In classical algorithms, subspace methods can transform the solution of high-dimensional linear systems into a series of lower-dimensional linear systems. We can apply subspace methods to QCFD algorithms to address scalability issues. However, subspace methods require high-precision solutions for each linear system, imposing a new challenge on QCFD algorithms, called high precision challenge.
@@
To address these challenges, we implement a comprehensive approach, as illustrated in Fig. 1, to ensure compatibility on NISQ devices and high-precision solutions. First, we introduce the Iterative-QLS approach for each small problem, enabling a high-precision quantum solver on noisy and limited-scale devices. The cycle is repeated until the error, a.k.a the residual of the approximated solution, is suppressed below an expected precision to meet the desired convergence condition demanded by the divide-and-conquer subspace method. Second, the less-noisy while limited-scale solutions for different Krylov subspaces are synthesized to derive the large-scale solution. To mitigate the quantum-classical information conversion bottleneck, a sparse tomography technique is developed to extract the desired information from the readout more efficiently.
@@ 3.2 Iterative-QLS: Improving the precision of noisy quantum linear solver via iterative scheme
To improve solution accuracy, an iterative method is utilized. In each loop, we first derive an approximate solution $\tilde{\boldsymbol{x}}_{i}$ by training VQLS. The training process of VQLS is similar to a standard VQLS process, where we exclude the Hadamard test and replace it with a loss function based on the results of tomography. B.3 shows how to evaluate the loss function and its gradient. The tomography process is the sparse tomography proposed in this paper, with details provided in Sec. 3.4.
@@
2:Give a initial solution $\tilde{\boldsymbol{x}}$.
@@
3.3 Scaling quantum linear solver to arbitrary scale via subspace method
To enhance scalability, we develop the subspace method inspired by the classical renowned Krylov subspace method and the generalized minimal residual (GMRES) method [39]. Instead of solving an $n$-dimensional linear system intractable large for near-term quantum computers, we iteratively build small spaces of lower dimension $m\ll n$ and find the approximate solution therein. More specific, we consider the $m$-order Krylov subspace $\mathcal{K}_{m}$ spanned by basis $\left\{\boldsymbol{b},A\boldsymbol{b},A^{2}\boldsymbol{b},...,A^{m}\boldsymbol{b}\right\}$.
This integration of iterative-QLS and subspace method is implemented via a quantum-classical hybrid procedure. As shown in Fig. 1, the subspace is constructed on classical computers, and a subspace linear system is generated and solved by quantum computers. The algorithm is presented in Alg. 2.
@@ 4.1 Experimental setup
To offer a comprehensive supplementary description of the computational process in this work, a specific detail has been added as follows:
@@
The experiments are conducted on a superconducting quantum computer, named “Wukong” which is fabricated by Origin Quantum. The detailed specification of Wukong is shown in A.
@@ 4.3 Acoustic wave propagation simulation
$\frac{\partial u}{\partial t}$ $=-\frac{1}{\bar{\rho}}\frac{\partial p}{\partial x}-\bar{u}\frac{\partial u}{\partial x},$
$\frac{\partial v}{\partial t}$ $=-\frac{1}{\bar{\rho}}\frac{\partial p}{\partial y}-\bar{u}\frac{\partial v}{\partial x}.$
@@
To accurately capture the sound waves, numerical schemes with low dissipation and low dispersion are required to discretize the equations. The dispersion relation preserving (DRP) scheme is one of the most efficient schemes widely applied in computational aeroacoustics (CAA). In this work, the seven-point DRP scheme is applied to spatial discretization. For time advance, the derivative is discretized with first order Euler implicit scheme to get higher advanced efficiency and more robust numerical stability. Furthermore, to avoid non-physical wave reflection near boundaries, the non-reflecting radiation boundary condition is imposed on all boundaries.
@@
As a result, a large sparse linear system of equations is obtained, of which the dimension is $3N$, where $N$ is the number of grid points. In practice, the grid point number is relatively large, so the linear system is too large for the current NISQ device to handle. Thus the Sub-VQLS can be used to reduce the problem size to match the capability of NISQ devices.
In this experiment, a 1000Hz sound source is placed at the center of a $[-1,1]\times[-1,1]$ square zone. The zone is uniformly meshed with $35\times 35$ grid points. In addition, three layers of ghost points are extended outside the boundaries for better numerical treatment near boundaries, resulting in actual $41\times 41$ grid points. To save time, the flow is advanced to a developed state with classical methods to get the initial flow field. The classical linear solver uses the LU decomposition method to yield results accurately enough. Starting with the initial flow field, the flow is solved with the subspace method on the quantum computer, and the wave propagates forward about 0.75 wavelengths. The final pressure field is compared with the classical counterpart to validate the reliability.
Starting with the same initial flow field, the flow is advanced with time step size. Fig. 4(a) and (b) show the pressure fluctuation contour of the classical and quantum results respectively. Both results are in qualitative agreement. For quantitative comparison, the results of slices at $Y=0.0$ are shown in Fig. 4(c). After a short evolution, a shift of acoustic waves propagating outward is observed, and the quantum result agrees with the classical result at phase and magnitude. Therefore, the proposed quantum approach is suitable for unsteady acoustic wave propagation and the simulation is reliable.
@@
Fig. 5(a) shows a piece of convergence history of the subspace method during simulation. For each time step, the residual converges after 3 subspace iterations, showing the good convergence property of Iterative-QLS method. To further investigate the convergence performance of Iterative-QLS, a piece of convergence history of the iteration of VQLS during simulation is shown in Fig. 5(b). It is observed that, in each subspace iteration of every time step, the residual of VQLS drops quickly during iteration, showing the fast convergence of solving the subspace linear system of equations with Iterative-QLS.
@@
5 Scalability of the proposed method
Sub-QLS represents a quantum-classical hybrid method that converts the original problem to a smaller-scale subproblem that can be executed on the current quantum computer. In this section, we will show that this idea has potential scalability: improving the utilized quantum computer in this algorithm will result in a corresponding increase in the algorithm performance.
In the way toward performing a practical scale of fluid simulation with the quantum computer, we are still facing a major limitation: the qubit’s quality and the qubit count. Firstly, we show that by decreasing the amount of qubit error, the total number of circuit executions can be significantly reduced. We conduct quantum circuit simulations for Poiseuille flow on a classical computer with different qubit errors. The qubit error is defined as the strength of the depolarizing error channel appended after every quantum gate. In Fig. 6, we can see that the count of the circuit execution significantly decreases with a decreasing error rate. Therefore, our method promises a faster execution with a better qubit quality that is utilized in the experiment.
Secondly, the qubit count in simulating the Poiseuille and acoustic wave propagation problems is only 2, mainly due to the imperfection of the quantum computer utilized in our experiments. We show that our method can be extended to four qubits, with additional results shown in E. Moreover, we argue that the integration of the subspace method on the flow simulation is promising when more qubits are provided to enable a larger subspace size $m=2^{n}$ where $n$ is the qubit and $m$ is the size of the subspace.
In Fig. 7, we solve a matrix of $1024$ dimension with the size of subspace varies from $n=4$ to $n=10$ where $n=\log_{2}m$ represents both the number of qubits and logarithm of the subspace. We plot the final residual, the number of steps of subspace, and predicted circuit execution time on a quantum computer with different qubit numbers. The results are divided into two regimes. At $n\leq 8$, where the subspace size is far less than the dimension, we call it the “NISQ regime”, representing the case when we have a very limited number of high-quality qubits. At $n\geq 9$, where the subspace size approaches the matrix size, we call it the “FTQC regime”, representing the case when we are able to put a full flow field on a quantum computer.
In the NISQ regime, we find the number of subspace steps follows a nearly exponential dependence on the number of qubits, while the system converges to a similar residual. This represents that the investment of every additional qubit will result in exponential computing power. The predicted execution time also follows the same trend, where additional costs will be incurred when we have to execute more gates.
In the FTQC regime, the subspace size approaches the dimension of the original matrix. As a result, the residual decreases significantly because it almost becomes a direct inversion algorithm. The number of subspace steps and the predicted execution time also decreases significantly when it enters such a regime. This represents that even in the FTQC regime, it is meaningful to explore the potential of integration of the subspace method and quantum computing before direct inversion is allowed.
Refer to caption NISQ Regime FTQC Regime
Figure 7: Results of solving a 1024-dimension matrix with different subspace sizes. (a) The residual of the final solution after convergence. (b) The number of subspace steps. (c) Predicted execution time on a quantum computer.

### v2

@@ Abstract
Quantum computational fluid dynamics (QCFD) offers a promising alternative to classical computational fluid dynamics (CFD) by leveraging quantum algorithms for higher efficiency. This paper introduces a comprehensive QCFD method implemented on a superconducting quantum computer, demonstrating successful simulations of steady Poiseuille flow and unsteady acoustic wave propagation. The Poiseuille flow simulation achieved a relative error of less than $0.2\%$, and the unsteady acoustic wave simulation solved a 5043-dimension matrix, marking the largest fluid simulation on a quantum computer to date. Our approach bridges quantum and classical computing, adapting to quantum hardware constraints and offering scalable solutions for large-scale CFD problems, which paves the way for practical applications of near-term quantum computers in computational science.
@@ 1 Introduction
Computational fluid dynamics (CFD) has played a pivotal role in numerous engineering and scientific disciplines, providing invaluable insights into complex flow behavior. However, as computational demands increase and the complexity of fluid dynamics problems intensifies, the limitations of classical computing architectures become increasingly apparent. Quantum computational fluid dynamics (QCFD) emerges as a promising alternative, offering the potential to overcome these limitations and unlock unprecedented computational power [1]. At its core, QCFD employs quantum algorithms and quantum computing hardware to simulate fluid dynamics phenomena with higher efficiency. The specific methods of QCFD are diverse, including Lagrangian-type methods [2, 3], explicit Euler-type methods [4], implicit Euler-type methods [5, 6], and methods based on quantum neural networks [7]. Focusing on solving complicated fluid governing equations, the linear system of equations plays a fundamental role in most classical CFD methods. While solving linear systems is the most time-consuming part of these methods, quantum linear solvers (QLS) promise an exponential speedup since first proposed in Ref. [8]. Tab. 1 reports these developments on QLS.
@@
Despite these prominent advancements, the implementation of QCFD still faces challenges based on near-term quantum devices, i.e., noisy intermediate-scale quantum (NISQ) devices [15]. These aforementioned QLS are designed in the context of the fault-tolerant quantum computation (FTQC) era [16, 17]. In a fault-tolerant quantum computer, redundant noisy physical qubits are employed to encode an ideal “logical qubit” without error [18]. However, only a few logical qubits can be demonstrated so far [19, 20], preventing most QLS from near-term practicability.
@@
Some efforts are conducted to enable computing fluid flows on quantum computers. One methodology is to emulate the flow field as a quantum state, utilizing the quantum state evolution to emulate the flow simulation, as shown in [21, 22]. Another methodology tries a variational approach to solve the equation, such as employing variational quantum linear solvers (VQLS) [23, 24, 25] in CFD, including solving potential and Stokes flows, lid-driven cavity, and the advection-diffusion equation [26, 27, 28, 29]. However, current implementations are limited to small sizes due to the inherent limitations of near-term quantum computers, which are significantly affected by noise. Additionally, encoding the entire flow field into a quantum state and reading it out into classical data using full quantum tomography is highly inefficient. These noise and readout issues impede the scalability of existing methods to larger problems.
@@
In this paper, we propose a comprehensive method that enables large-scale and high-precision fluid simulation on near-term quantum computers. We implement it on a superconducting quantum computer, successfully simulating steady Poiseuille flow and unsteady acoustic wave propagation to showcase the ability of our method. In the Poiseuille flow simulation, we achieve less than $0.2\%$ relative error for showcasing the accuracy of a single step. In the unsteady acoustic wave propagation simulation, we successfully simulated 1681 grid points. Each time step solves a 5043-dimension matrix, which is the largest fluid simulation on the quantum computer reported so far.
@@
To mitigate experimental error arising from noisy devices and approximation error, we introduce Iterative-QLS, by introducing iterations to enhance the precision rather than employing a single VQLS. This cycle aims to diminish residual error by achieving synthetic accuracy levels even lower than the physical accuracy. To overcome the problem size constraints imposed by intermediate-scale quantum processors, the subspace method is employed. This approach enables the hybrid utilization of quantum computing and classical computing, enabling simulation on a larger flow field. Furthermore, to enhance the efficiency of extracting classical information from quantum states, we have developed the sparse tomography technique. In summary, we present an integrated QCFD protocol capable of deriving high-precision solutions for problems of arbitrary size.
@@
We regard our algorithm as a bridge between quantum and classical computers. According to our findings, our method can adapt to the quantum computer’s capabilities, enabling collaboration between quantum and classical computers. Although the current solution uses only a limited number of qubits due to hardware constraints, advancements in quantum computing offer the possibility of solving larger subspaces by investing more quantum resources. Our approach is also a general-purpose quantum-enhanced linear solver method, applicable to other potential applications in computational science, including but not limited to finite element analysis, computer graphics, and signal processing. With successful, high-precision demonstrations on large-scale CFD problems, this paper marks a step towards scalable quantum-enhanced solutions for CFD and paves the way for the practical application of near-term quantum computers.
@@ 2.1 Basic concepts in quantum computing
The basic information unit in quantum information is the qubit, which is analogous to a classical bit but with the ability to exist in a superposition of states. Mathematically, a single qubit state can be represented as a vector in 2-dimension Hilbert space:
@@ 2.2 The technical path of QCFD using QLS
In CFD, one must solve linear or nonlinear partial differential equations, such as the Navier-Stokes equations. These equations can be represented as
$\partial{Q}/\partial{t}=\mathcal{F}(Q),$ (1)
where $\mathcal{F}(Q)$ represents linear or nonlinear operators. The first step is discretization, which involves breaking down the equations into a form that can be handled computationally. Both time and space discretization schemes are employed, which can be either explicit or implicit. While explicit methods have strict stability requirements limiting the time step, implicit methods are preferred for their ability to handle larger time steps and provide numerical stability, particularly for stiff differential equations. Therefore, we use implicit methods, which are formulated as
@@
$\frac{\vec{Q}^{t+1}-\vec{Q}^{t}}{\Delta}=\vec{R}(\vec{Q}^{t+1}),$ (3)
where $\vec{Q}^{t+1}\in\mathbb{R}^{n}$, and $n$ represents the dimension of the discretized space. Generally, when $\mathcal{F}(Q)$ is nonlinear, this leads to a nonlinear system of equations for $\vec{Q}^{t+1}$, necessitating a linearization procedure.
@@ 2.3 Challenges of implementing CFD simulation on NISQ devices
Currently, several QCFD algorithms have been proposed, but there are still challenges when solving specific CFD problems on NSIQ devices. Here, we list and analyze these challenges.
Firstly, due to the impact of quantum noise, the quantum state produced by a quantum circuit executed on NISQ devices is prone to errors. This inaccuracy results in low-precision, unreliable solutions and subsequently hinders the computation of unsteady flow using time-stepping methods.
Secondly, the number of qubits on NISQ devices is limited. This limitation arises not only because the number of qubits that can be fabricated on a quantum processor is restricted, but also because quantum noise limits the size of the circuit, preventing the number of qubits used in a quantum circuit from increasing [37].
@@ 3.1 Overview
When solving large-scale problems, the number of qubits required by QCFD algorithms may exceed the chip’s capacity. Therefore, a key issue for QCFD algorithms is how to solve large-scale CFD problems on NISQ devices with a limited number of qubits. In classical algorithms, subspace methods can transform the solution of high-dimensional linear systems into a series of lower-dimensional linear systems. We can apply subspace methods to QCFD algorithms to address scalability issues. However, subspace methods require high-precision solutions for each linear system, imposing a new challenge on QCFD algorithms, called high precision challenge.
@@
To address these challenges, we implement a comprehensive approach, as illustrated in Fig. 1, to ensure compatibility on NISQ devices and high-precision solutions. First, we introduce the Iterative-QLS approach for each small problem, enabling a high-precision quantum solver on noisy and limited-scale devices. The cycle is repeated until the residual of the approximated solution is suppressed below an expected precision to meet the desired convergence condition demanded by the divide-and-conquer subspace method. Second, the less-noisy while limited-scale solutions for different Krylov subspaces are synthesized to derive the large-scale solution. To mitigate the quantum-classical information conversion bottleneck, a sparse tomography technique is developed to extract the desired information from the readout more efficiently.
@@ 3.2 Iterative-QLS: Improving the precision of noisy quantum linear solver via iterative scheme
To improve solution accuracy, an iterative method is utilized. In each loop, we first derive an approximate solution $\tilde{\boldsymbol{x}}_{i}$ by training VQLS. The training process of VQLS is similar to a standard VQLS process, where we exclude the Hadamard-Test and replace it with a loss function based on the results of tomography. B.3 shows how to evaluate the loss function and its gradient. The tomography process is the sparse tomography proposed in this paper, with details provided in Sec. 3.4.
@@
2:Give an initial solution $\tilde{\boldsymbol{x}}$.
@@
3.3 Scaling quantum linear solver to arbitrary scale via the subspace method
To enhance scalability, we develop the subspace method inspired by the classical renowned Krylov subspace method and the generalized minimal residual (GMRES) method [41]. Instead of solving an $n$-dimensional linear system intractable large for near-term quantum computers, we iteratively build small spaces of lower dimension $m\ll n$ and find the approximate solution therein. More specific, we consider the $m$-order Krylov subspace $\mathcal{K}_{m}$ spanned by basis $\left\{\boldsymbol{b},A\boldsymbol{b},A^{2}\boldsymbol{b},...,A^{m}\boldsymbol{b}\right\}$. This integration of Iterative-QLS and subspace method is implemented via a quantum-classical hybrid procedure. As shown in Fig. 1, the subspace is constructed on classical computers, and a subspace linear system is generated and solved by quantum computers. The algorithm is presented in Alg. 2.
@@ 3.3 Scaling quantum linear solver to arbitrary scale via the subspace method
Iterative-QLS is chosen as the internal solver for the subspace method primarily due to its high-precision features. Replacing Iterative-QLS with a single VQLS process may cause divergence in the subspace process, as each integration is not guaranteed to sufficiently shrink the residual, considering the error produced in the real quantum computer. Furthermore, Iterative-QLS can be replaced by a faster and more accurate quantum solver, which remains a subject for future work.
The size of the subspace can be determined by the capability of the internal QLS process. By providing a more powerful quantum computer, one can utilize a larger subspace size to accelerate the iteration process, an idea widely discussed in the literature [42, 43, 44, 45, 46, 47]. Since QLS can solve $2^{n}$ linear system with $\operatorname{poly}(n)$ qubits, it offers a broad potential for quantum computing to accelerate at least one part of the fluid simulation. We conduct simulations on examples emulating scenarios with stronger quantum computers, detailed in F.
@@ 4.1 Experimental setup
To offer a comprehensive supplementary description of the computational process in this work, details of QLS have been added as follows:
@@
The experiments are conducted on a superconducting quantum computer, named “Wukong” which is fabricated by Origin Quantum. The detailed specification of Wukong is shown in A. No additional quantum error mitigation technique [52] is employed in experiments.
@@ 4.3 Acoustic wave propagation simulation
$\frac{\partial u}{\partial t}$ $=-\frac{1}{\bar{\rho}}\frac{\partial p}{\partial x}-\bar{u}\frac{\partial u}{\partial x}-\bar{v}\frac{\partial u}{\partial y},$
$\frac{\partial v}{\partial t}$ $=-\frac{1}{\bar{\rho}}\frac{\partial p}{\partial y}-\bar{u}\frac{\partial v}{\partial x}-\bar{v}\frac{\partial v}{\partial y}.$
@@
To accurately capture the sound waves, numerical schemes with low dissipation and low dispersion are required to discretize the equations. The dispersion relation preserving (DRP) scheme [53] is one of the most efficient schemes widely applied in computational aeroacoustics (CAA). In this work, the seven-point DRP scheme [54] is applied to spatial discretization. For time advance, the derivative is discretized with first order Euler implicit scheme to get higher advanced efficiency and more robust numerical stability. Furthermore, to avoid non-physical wave reflection near boundaries, the non-reflecting radiation boundary condition is imposed on all boundaries [55].
@@
As a result, a large sparse linear system of equations is obtained, of which the dimension is $3N$, where $N$ is the number of grid points. In practice, the grid point number is relatively large, so the linear system is too large for the current NISQ device to handle. Thus the subspace method can be used to reduce the problem size to match the capability of NISQ devices.
In this experiment, a 1000Hz sound source is placed at the center of a $[-1,1]\times[-1,1]$ square zone. The zone is uniformly meshed with $35\times 35$ grid points. In addition, three layers of ghost points are extended outside the boundaries for better numerical treatment near boundaries, resulting in actual $41\times 41$ grid points. To save time, the flow is advanced to a developed state with classical methods to get the initial flow field. The classical linear solver uses the LU decomposition method to yield results accurately enough. Starting with the initial flow field, the flow is solved with the subspace method on the quantum computer, and the wave propagates forward about $0.93$ wavelengths. The final pressure field is compared with the classical counterpart to validate the reliability.
Starting with the same initial flow field, the flow is advanced with time step size. Fig. 4(a) and (b) show the pressure fluctuation contour of the quantum and classical results respectively. Both results are in qualitative agreement. For quantitative comparison, the results of slices at $Y=0.0$ are shown in Fig. 4(c). After a short evolution, a shift of acoustic waves propagating outward is observed, and the quantum result agrees with the classical result at phase and magnitude. Therefore, the proposed quantum approach is suitable for unsteady acoustic wave propagation and the simulation is reliable.
@@
Fig. 5(a) shows a piece of convergence history of the subspace method during simulation. For each time step, the residual converges after 3 subspace iterations, showing the good convergence property of Iterative-QLS method. To further investigate the convergence performance of Iterative-QLS, a piece of convergence history of the iteration of VQLS during simulation is shown in Fig. 5(b). It is observed that, in each subspace iteration of every time step, the residual of Iterative-QLS drops quickly during iteration, showing the fast convergence of solving the linear system of equations with the subspace method.